# Resolving phonon-mediated superconducting pairing symmetries from first-principles calculation

### v2

@@ Abstract
The quest for topological superconductors triggers revived interests in resolving non-$s$-wave pairing channels mediated by phonons. While density functional theory and denstify functional perturbtaion theory have established a powerful framework to calculate electron-phonon couplings in real materials in a first-principles way, its application is largely limited to conventional $s$-wave superconductivity. Here, we formulate an efficient and simple-to-use algorithm for first-principles pairing channel analysis, and apply it to several representative material systems.
@@ I Introduction
Recently, the potential role of phonons in achieving unconventional superconductivity attracts revived theoretical attentions. The concern is that while electron-electron Coulomb repulsion always suppresses the conventional $s$-wave pairing, it can be collaborative with phonons in stabilizing other pairing symmetries, e.g. in a topological superconductor [5, 6, 7], or in high-Tc cuprate and pnictides superconductors [8, 9, 10]. In these cases, the subleading non-$s$-wave pairing channels from EPC should be taken into account.
@@
Nomoto et al. [12] extended DFT for superconductivity (SCDFT) [13, 14] to deal with odd-parity pairing in an ab initio way. The treatment is exact when $S_{z}$ is a good quantum number, e.g. without the spin-orbit coupling (SOC). When SOC is strong, however, the exact forms of odd-parity pairing in the basis of Kohn-Sham (K-S) eigenstates are not known a priori.
The primary goal of this Article is to formulate an efficient and simple-to-use algorithm for first-principles EPC pairing-channel analysis. The key is to properly solve gauge indeterminancy of the EPCs. After that, we show that the pairing channels can be resolved by standard matrix diagonalization. We call this approach the “direct diagonalization method” (DDM). In Sec. II, we will first review the general theory of phonon-mediated pairing interactions, and then discuss the gauge complexity. In the end, we present the DDM method. In Sec. III, the DDM method is applied to several realistic material systems. Section IV concludes this Article.
@@ II.1 General theory
We start by considering a system with inversion ($\mathcal{I}$) and time-reversal ($\mathcal{T}$) symmetries. Without $\mathcal{I}$, the definition of parity is missing, and the gap function is in general a mixture of odd and even parities. Without $\mathcal{T}$, the scenario is complicated by the coexistence of magnetism, which we defer for future works. The combination of $\mathcal{I}$ and $\mathcal{T}$ enforces that every eigenstate is at least doubly degenerate. Every K-S eigenstates can then be labeled by a band index $n$, the lattice momentum $\mathbf{k}$ and a pseudospin variable $s=\pm$.
@@
$H_{ep}=\sum_{n,m,\mathbf{k,q},s,s^{\prime}\upsilon}g_{n\mathbf{k}s,m\mathbf{k+q}s^{\prime}}^{\upsilon}c_{m\mathbf{k+q}s^{\prime}}^{\dagger}c_{n\mathbf{k}s}(b_{\upsilon,\mathbf{q}}^{\dagger}+b_{\upsilon,-\mathbf{q}}),$
in which $c_{n\mathbf{k}s}$ ($b_{\upsilon,q}$) is the annihilation operator of the electron (phonon) eigenstate. The phonons are labeled by a mode index $\upsilon$ and the lattice vection $\mathbf{q}$. $g_{n\mathbf{k}s,m\mathbf{k+q}s^{\prime}}^{\upsilon}$ is the EPC coefficient.
@@
$H_{pair}=-\sum_{n\mathbf{k},m\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ (1)
$c_{m\mathbf{k}^{\prime}s_{3}}^{\dagger}c_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}c_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$
@@
$V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}\frac{g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}g_{n\mathbf{-k}s_{2},m\mathbf{-k^{\prime}}s_{4}}^{\upsilon}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (2)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}),$
in which $\epsilon_{n\mathbf{k}}$ ($\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}$) is the electron (phonon) eigen-energy, and we restrict the involved electrons residing at the Fermi energy ($\epsilon_{F}$).
To bridge the interaction to a matrix analysis problem, it is convenient to view $V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ as the components of a matrix $\mathbb{V}$ and define a vector operator $\hat{\Delta}$ with its components
$\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=c_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$ (3)
which physically annihilates a Copper pair consisting of electrons at the $|n\mathbf{k}s_{1}\rangle$ and $|n,\mathbf{-k},s_{2}\rangle$ states. Its ground-state expectation value $\langle\hat{\Delta}(n\mathbf{k}s_{1}s_{2})\rangle$ reflects the SC order parameter. Eq. (1) is then transformed into a compact quadratic form:
$H_{pair}=-\hat{\Delta}^{\dagger}\mathbb{V}\hat{\Delta}.$ (4)
@@ II.2 Gauge complexity
The complexity is that EPCs are routinely evaluated on a $k$-mesh in the full Brillouin zone, where K-S orbitals at each $k$ point is treated independently with “gauge” indeterminacy. It is straightforward to see that $\hat{\Delta}(n\mathbf{k}s_{1}s_{2})$ is gauge dependent by adding an arbitrary U(1) phase to $c_{n\mathbf{k}s_{1}}$. More critically, the double degeneracy allows any recombination of the degenerate K-S orbitals as the eigenstate. Therefore, an arbitrary SU(2) rotation usually also presents in the first-principles data, especially for strong SOC systems, which hinders a textbook definition of the pairing symmetry assuming $s_{1},s_{2}$ in $\hat{\Delta}$ refer to a global spin axis. It is worth mentioning that maximally localized Wannier function interpolation is a powerful procedure for ab initio gauge fixing [15], but to properly specify the pseudo spin index, especially for strong SOC systems, could still be tricky.
@@ II.3 DDM solution
The main advantage of the DDM method is to avoid specifying the gauge condition, such that the computational procedure could label the pairing channel in any gauge. Let us denote
$\mathcal{T}c_{n\mathbf{k}s}\equiv s\tilde{c}_{n,-\mathbf{k},-s},$ (5)
in which the sign $s$ before $\tilde{c}$ is used to explicitly track the $\mathcal{T}^{2}=-1$ condition for an electron. The tilde symbol implies that $(\tilde{c}_{n,-\mathbf{k},+},\tilde{c}_{n,-\mathbf{k},-})$ is allowed to differ from $(c_{n,-\mathbf{k},+},c_{n,-\mathbf{k},-})$, i.e. the numerical K-S states at $-\mathbf{k}$ without gauge fixing, by an arbitrary U(1)$\times$×SU(2) transformation. Thus defined $\tilde{c}_{n,-\mathbf{k},-s}$ also relates to $c_{n\mathbf{k},-s}$ via spatial inversion as:
$\mathcal{I}c_{n\mathbf{k},-s}=e^{i\alpha_{n\mathbf{k}}}\tilde{c}_{n,-\mathbf{k},-s}.$ (6)
Note that theoretical analysis usually implicitly assumes $e^{i\alpha_{n\mathbf{k}}}=1$, which however does not automatically hold for the numerical K-S states. A U(1) transformation $c_{n\mathbf{k}\pm}\rightarrow e^{i\alpha_{n\mathbf{k}\pm}}c_{n\mathbf{k}\pm}$ will lead to an overall phase shift $\alpha_{n\mathbf{k}}=\alpha_{n\mathbf{k}+}+\alpha_{n\mathbf{k}-}$.
We rewrite Eqs. (1) and (2) into
$H_{pair}=-\sum_{n\mathbf{k},m\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ (7)
$c_{m\mathbf{k}^{\prime}s_{3}}^{\dagger}\tilde{c}_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}},$
@@
$V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}s_{2}s_{4}\frac{g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}(g_{n\mathbf{k},-s_{2},m\mathbf{k^{\prime}},-s_{4}}^{\upsilon})^{*}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (8)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}).$
The emergence of summand sign $s_{2}s_{4}$ is a direct consequence of $\mathcal{T}^{2}=-1$ [c.f. Eq. (5)]. A useful point is that the EPC between $\tilde{c}_{m,\mathbf{-k}^{\prime},s^{\prime}}^{\dagger}$ and $\tilde{c}_{n,\mathbf{-k},s}$ can be readily obtained from that between $c_{m,\mathbf{k}^{\prime},-s^{\prime}}^{\dagger}$ and $c_{n,\mathbf{k},-s}$ by applying $\mathcal{T}$, so the tilde states never need to be explicitly constructed.
@@
$\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}}.$ (9)
@@
$\mathcal{I}\hat{\Delta}(n\mathbf{k}\pm\pm)=-\hat{\Delta}(n\mathbf{k}\pm\pm)$ (10)
$\mathcal{I}\hat{\Delta}(n\mathbf{k}\pm\mp)=-\hat{\Delta}(n\mathbf{k}\mp\pm).$
Note that the undertermined U(1) phase $e^{i\alpha_{n\mathbf{k}}}$ cancels.
@@
$\hat{\Delta}^{S}(n\mathbf{k})=\frac{1}{2}[\hat{\Delta}(n\mathbf{k}+-)-\hat{\Delta}(n\mathbf{k}-+)],$ (11)
@@
$\hat{\Delta}^{T_{0}}(n\mathbf{k})$ $=$ $\frac{1}{2}[\hat{\Delta}(n\mathbf{k}+-)+\hat{\Delta}(n\mathbf{k}-+)]$
$\hat{\Delta}^{T_{\pm}}(n\mathbf{k})$ $=$ $\hat{\Delta}(n\mathbf{k}\pm\pm).$ (12)
Transforming into this basis, the kernel $\mathbb{V}$ is block diagonalized into the even-parity singlet sector ($\mathbb{V}^{S}$) and the the odd-parity triplet sector ($\mathbb{V}^{T}$). The matrix elements are:
$V^{S}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon s_{1}s_{3}}\frac{|g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}|^{2}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (13)
$\times$ $\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F}),$
@@
$V^{T}(n\mathbf{k},m\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon}\begin{pmatrix}g_{n\mathbf{k}+,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&-\sum_{s}g_{n\mathbf{k}+,m\mathbf{k^{\prime}}s}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}s}^{\upsilon})^{*}&-g_{n\mathbf{k}+,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}-,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\\ -\sum_{s}g_{n\mathbf{k}s,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k}s,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&\sum_{ss^{\prime}}ss^{\prime}|g_{n\mathbf{k}s_{1},m\mathbf{k^{\prime}}s_{3}}^{\upsilon}|^{2}&-\sum_{s}g_{n\mathbf{k}s,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}s,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\\ -g_{n\mathbf{k}-,m\mathbf{k^{\prime}}+}^{\upsilon}(g_{n\mathbf{k},+,m\mathbf{k^{\prime}}-}^{\upsilon})^{*}&-\sum_{s}g_{n\mathbf{k}-,m\mathbf{k^{\prime}}s}^{\upsilon}(g_{n\mathbf{k}+,m\mathbf{k^{\prime}}s}^{\upsilon})^{*}&g_{n\mathbf{k}-,m\mathbf{k^{\prime}}-}^{\upsilon}(g_{n\mathbf{k}+,m\mathbf{k^{\prime}}+}^{\upsilon})^{*}\end{pmatrix}$ (14)
$\times$ $\frac{\delta(\epsilon_{n\mathbf{k}}-\epsilon_{F})\Theta(\epsilon_{m\mathbf{k}\textquoteright}-\epsilon_{F})}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}.$
It is heuristic to inspect that the average of $V^{S}(n\mathbf{k},m\mathbf{k}^{\prime})$ normalized by the Fermi surface density of states is the isotropic dimensionless EPC strength $\lambda$ commonly used to predict the $s$-wave Tc [1], which implicitly assumes that the matrix components do not contain any singular variation. The weak-SOC approximation employed by Nomoto et al. [12] to evaluate odd-parity pairing can be viewed as neglecting the off-diagonal matrix elements in Eq. (14).
@@
In practice, we non-dimensionlize $\mathbb{V}$ by $V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})\rightarrow V_{s_{2},s_{1};s_{3},s_{4}}(n\mathbf{k},m\mathbf{k}^{\prime})N_{n\mathbf{k}}$, where $N_{n\mathbf{k}}$ is the contribution of the $|n\mathbf{k}s\rangle$ state to the total density of states. This makes the largest $\lambda^{S}_{i}$, which we denote as $\lambda^{S}_{max}$, directly comparable to the isotropic dimensionless EPC strength $\lambda$ reported in literature. The largest $\lambda^{T}_{i}$ ($\lambda^{T}_{max}$) gives the leading odd-parity spin-triplet pairing strength.

### v3

@@ Abstract
The quest for topological superconductors triggers revived interests in resolving non-$s$-wave pairing channels mediated by phonons. While density functional theory and density functional perturbation theory have established a powerful framework to calculate electron-phonon couplings in real materials in a first-principles way, its application is largely limited to conventional $s$-wave superconductivity. Here, we formulate an efficient and simple-to-use algorithm for first-principles pairing channel analysis, and apply it to several representative material systems.
@@ I Introduction
Recently, the potential role of phonons in achieving unconventional superconductivity attracts revived theoretical attentions. The concern is that while electron-electron Coulomb repulsion always suppresses the conventional $s$-wave pairing, it can be collaborative with phonons in stabilizing other pairing symmetries, e.g., in a topological superconductor [5, 6, 7], or in high-$T_{c}$ cuprate and pnictides superconductors [8, 9, 10]. In these cases, the non-$s$-wave pairing channels from EPC should be taken into account.
@@
Nomoto $et$ $al.$ [12] extended DFT for superconductivity (SCDFT) [13, 14] to deal with odd-parity pairing in an ab initio way. The treatment is exact when $S_{z}$ is a good quantum number, e.g., without the spin-orbit coupling (SOC). When SOC is strong, however, the exact forms of odd-parity pairing in the basis of Kohn-Sham (KS) eigenstates are not known a priori.
The primary goal of this article is to formulate an efficient and simple-to-use algorithm for first-principles EPC pairing-channel analysis. The key is to properly solve gauge indeterminacy of the EPCs. After that, we show that the pairing channels can be resolved by standard matrix diagonalization. We call this approach the “direct diagonalization method” (DDM). In Sec. II, we will first review the general theory of phonon-mediated pairing interactions, and then discuss the gauge complexity. In the end, we present the DDM. In Sec. III, the DDM is applied to several realistic material systems. Section IV concludes this article.
@@ II.1 General theory
We start by considering a system with inversion ($\mathcal{I}$) and time-reversal ($\mathcal{T}$) symmetries. Without $\mathcal{I}$, the definition of parity is missing, and the gap function is in general a mixture of odd and even parities. Without $\mathcal{T}$, the scenario is complicated by the coexistence of magnetism, which we defer for future works. The combination of $\mathcal{I}$ and $\mathcal{T}$ enforces that every eigenstate is at least doubly degenerate. Every KS eigenstate can then be labeled by a band index $n$, the lattice momentum $\mathbf{k}$ and a pseudospin variable $s=\pm$.
@@
$H_{ep}=\sum_{n,m,\mathbf{k,q},s,s^{\prime}\upsilon}g_{n,\mathbf{k},s;m,\mathbf{k+q},s^{\prime}}^{\upsilon}c_{m,\mathbf{k+q},s^{\prime}}^{\dagger}c_{n,\mathbf{k},s}(b_{\upsilon,-\mathbf{q}}^{\dagger}+b_{\upsilon,\mathbf{q}}),$
in which $c_{n,\mathbf{k},s}$ ($b_{\upsilon,\mathbf{q}}$) is the annihilation operator of the electron (phonon) eigenstate. The phonons are labeled by a mode index $\upsilon$ and the lattice vector $\mathbf{q}$. $g_{n,\mathbf{k},s;m,\mathbf{k+q},s^{\prime}}^{\upsilon}$ is the EPC coefficient.
@@
$H_{pair}=-\frac{1}{2}\sum_{n,\mathbf{k},m,\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ (1)
$c_{m,\mathbf{k}^{\prime},s_{3}}^{\dagger}c_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}c_{n,\mathbf{-k},s_{2}}c_{n,\mathbf{k},s_{1}},$
@@
$V_{s_{1},s_{2},s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ $=$ $2\sum_{\upsilon}\frac{g_{n,\mathbf{k},s_{1};m,\mathbf{k^{\prime}},s_{3}}^{\upsilon}g_{n,\mathbf{-k},s_{2};m,\mathbf{-k^{\prime}},s_{4}}^{\upsilon}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (2)
$\times$ $\delta(\epsilon_{n,\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m,\mathbf{k}\textquoteright}-\epsilon_{F}),$
in which $\epsilon_{n,\mathbf{k}}$ ($\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}$) is the electron (phonon) eigen-energy, and we restrict the involved electrons residing at the Fermi energy ($\epsilon_{F}$). We do not consider strong coupling systems beyond this treatment.
To bridge the interaction to a matrix analysis problem, it is convenient to view $V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ as the components of a matrix $\mathbb{V}$ and define a vector operator $\hat{\Delta}$ with its components
$\hat{\Delta}(n,\mathbf{k},s_{1},s_{2})=c_{n,\mathbf{-k},s_{2}}c_{n,\mathbf{k},s_{1}},$ (3)
which physically annihilates a Copper pair consisting of electrons at the $|n,\mathbf{k},s_{1}\rangle$ and $|n,\mathbf{-k},s_{2}\rangle$ states. Its ground-state expectation value $\langle\hat{\Delta}(n,\mathbf{k},s_{1},s_{2})\rangle$ reflects the SC order parameter. Eq. (1) is then transformed into a compact quadratic form:
$H_{pair}=-\frac{1}{2}\hat{\Delta}^{\dagger}\mathbb{V}\hat{\Delta}.$ (4)
@@ II.2 Gauge complexity
The complexity is that EPCs are routinely evaluated on a $k$-mesh in the full Brillouin zone, where KS orbitals at each $k$ point are treated independently with ”gauge” indeterminacy. It is straightforward to see that $\hat{\Delta}(n,\mathbf{k},s_{1},s_{2})$ is gauge dependent by adding an arbitrary U(1) phase to $c_{n,\mathbf{k},s_{1}}$. More critically, the double degeneracy allows any recombination of the degenerate KS orbitals as the eigenstate. Therefore, an arbitrary SU(2) rotation is also presented in the first-principles data, especially for strong SOC systems, which hinders a textbook definition of the pairing symmetry assuming $s_{1},s_{2}$ in $\hat{\Delta}$ refer to a global spin axis. It is worth mentioning that maximally localized Wannier function interpolation is a powerful procedure for ab initio gauge fixing [15], but to properly specify the pseudospin index, especially for strong SOC systems, could still be tricky.
@@ II.3 DDM solution
The main advantage of the DDM is to avoid specifying the gauge condition, such that the computational procedure could label the pairing channel in any gauge. Let us denote
$\mathcal{T}c_{n,\mathbf{k},s}\equiv s\tilde{c}_{n,-\mathbf{k},-s},$ (5)
in which the sign $s$ before $\tilde{c}$ is used to explicitly track the $\mathcal{T}^{2}=-1$ condition for an electron. The tilde symbol implies that $(\tilde{c}_{n,-\mathbf{k},+},\tilde{c}_{n,-\mathbf{k},-})$ is allowed to differ from $(c_{n,-\mathbf{k},+},c_{n,-\mathbf{k},-})$, i.e., the numerical KS states at $-\mathbf{k}$ without gauge fixing, by an arbitrary U(1)$\times$×SU(2) transformation. Thus defined $\tilde{c}_{n,-\mathbf{k},-s}$ also relates to $c_{n,\mathbf{k},-s}$ via spatial inversion as:
$\mathcal{I}c_{n,\mathbf{k},-s}=e^{i\alpha_{n,\mathbf{k}}}\tilde{c}_{n,-\mathbf{k},-s}.$ (6)
Note that theoretical analysis usually implicitly assumes $e^{i\alpha_{n,\mathbf{k}}}=1$, which however does not automatically hold for the numerical KS states. A U(1) transformation $c_{n,\mathbf{k},\pm}\rightarrow e^{i\alpha_{n,\mathbf{k},\pm}}c_{n,\mathbf{k},\pm}$ will lead to an overall phase shift $\alpha_{n,\mathbf{k}}=\alpha_{n,\mathbf{k},+}+\alpha_{n,\mathbf{k},-}$ in Eq. (6).
By switching to the $\tilde{c}$ operators for all the -$\mathbf{k}$ states, we rewrite Eqs. (1) and (2) into
$H_{pair}=-\frac{1}{2}\sum_{n,\mathbf{k},m,\mathbf{k}^{\prime},s_{1,2,3,4}}V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ (7)
$c_{m,\mathbf{k}^{\prime},s_{3}}^{\dagger}\tilde{c}_{m,\mathbf{-k}^{\prime},s_{4}}^{\dagger}\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n,\mathbf{k},s_{1}},$
@@
$V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})=$
$2\sum_{\upsilon}s_{2}$ $s_{4}\frac{g_{n,\mathbf{k},s_{1};m,\mathbf{k^{\prime}},s_{3}}^{\upsilon}(g_{n,\mathbf{k},-s_{2};m,\mathbf{k^{\prime}},-s_{4}}^{\upsilon})^{*}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (8)
×$\times\delta(\epsilon_{n,\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m,\mathbf{k}\textquoteright}-\epsilon_{F}).$
The emergence of summand sign $s_{2},s_{4}$ is a direct consequence of $\mathcal{T}^{2}=-1$ [cf. Eq. (5)]. A useful point is that the EPC between $\tilde{c}_{m,\mathbf{-k}^{\prime},s^{\prime}}^{\dagger}$ and $\tilde{c}_{n,\mathbf{-k},s}$ can be readily obtained from that between $c_{m,\mathbf{k}^{\prime},-s^{\prime}}^{\dagger}$ and $c_{n,\mathbf{k},-s}$ by applying $\mathcal{T}$, so the tilde states never need to be explicitly constructed.
@@
$\hat{\Delta}(n,\mathbf{k},s_{1},s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n,\mathbf{k},s_{1}}.$ (9)
@@
$\mathcal{I}\hat{\Delta}(n,\mathbf{k},s_{2},s_{1})=-\hat{\Delta}(n,\mathbf{k},s_{1},s_{2}).$ (10)
Note that the undetermined U(1) phase $e^{i\alpha_{n,\mathbf{k}}}$ cancels.
@@
$\hat{\Delta}^{S}(n,\mathbf{k})=\frac{1}{\sqrt{2}}[\hat{\Delta}(n,\mathbf{k},+,-)-\hat{\Delta}(n,\mathbf{k},-,+)],$ (11)
@@
$\hat{\Delta}^{T_{0}}(n,\mathbf{k})$ $=$ $\frac{1}{\sqrt{2}}[\hat{\Delta}(n,\mathbf{k},+,-)+\hat{\Delta}(n,\mathbf{k},-,+)]$
$\hat{\Delta}^{T_{\pm}}(n,\mathbf{k})$ $=$ $\hat{\Delta}(n,\mathbf{k},\pm,\pm).$ (12)
Transforming into this basis, the kernel $\mathbb{V}$ is block diagonalized into the even-parity singlet sector ($\mathbb{V}^{S}$) and the odd-parity triplet sector ($\mathbb{V}^{T}$). The matrix elements are:
$V^{S}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ $=$ $\sum_{\upsilon,s_{1},s_{3}}\frac{|g_{n,\mathbf{k},s_{1};m,\mathbf{k^{\prime}},s_{3}}^{\upsilon}|^{2}}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}$ (13)
$\times$ $\delta(\epsilon_{n,\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m,\mathbf{k}\textquoteright}-\epsilon_{F}),$
@@
$V^{T}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ $=$ $2\sum_{\upsilon}\begin{pmatrix}g_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},+}^{\upsilon}(g_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},-}^{\upsilon})^{*}&-\sum_{s}\frac{sg_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},s}^{\upsilon}(g_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},s}^{\upsilon})^{*}}{\sqrt{2}}&-g_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},-}^{\upsilon}(g_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},+}^{\upsilon})^{*}\\ -\sum_{s}\frac{sg_{n,\mathbf{k},s;m,\mathbf{k^{\prime}},+}^{\upsilon}(g_{n,\mathbf{k},s;m,\mathbf{k^{\prime}},-}^{\upsilon})^{*}}{\sqrt{2}}&\sum_{ss^{\prime}}\frac{ss^{\prime}|g_{n,\mathbf{k},s;m,\mathbf{k^{\prime}},s^{\prime}}^{\upsilon}|^{2}}{2}&\sum_{s}\frac{sg_{n,\mathbf{k},s;m,\mathbf{k^{\prime}},-}^{\upsilon}(g_{n,\mathbf{k},s;m,\mathbf{k^{\prime}},+}^{\upsilon})^{*}}{\sqrt{2}}\\ -g_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},+}^{\upsilon}(g_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},-}^{\upsilon})^{*}&\sum_{s}\frac{sg_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},s}^{\upsilon}(g_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},s}^{\upsilon})^{*}}{\sqrt{2}}&g_{n,\mathbf{k},-;m,\mathbf{k^{\prime}},-}^{\upsilon}(g_{n,\mathbf{k},+;m,\mathbf{k^{\prime}},+}^{\upsilon})^{*}\end{pmatrix}$ (14)
$\times$ $\frac{\delta(\epsilon_{n,\mathbf{k}}-\epsilon_{F})\delta(\epsilon_{m,\mathbf{k}\textquoteright}-\epsilon_{F})}{\omega_{\mathbf{k^{\prime}}-\mathbf{k},\upsilon}}.$
It is heuristic to inspect that the average of $V^{S}(n,\mathbf{k};m,\mathbf{k}^{\prime})$ normalized by the Fermi surface density of states is the isotropic dimensionless EPC strength $\lambda$ commonly used to predict the $s$-wave $T_{c}$ [1], which implicitly assumes that the matrix components do not contain any singular variation. The weak-SOC approximation employed by Nomoto $et$ $al.$ [12] to evaluate odd-parity pairing can be viewed as neglecting the off-diagonal matrix elements in Eq. (14).
@@
In practice, we non-dimensionalize $\mathbb{V}$ by $V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})\rightarrow V_{s_{1},s_{2};s_{3},s_{4}}(n,\mathbf{k};m,\mathbf{k}^{\prime})N_{n,\mathbf{k}}$, where $N_{n,\mathbf{k}}$ is the contribution of the $|n,\mathbf{k},s\rangle$ state to the total density of states. This makes the largest $\lambda^{S}_{i}$, which we denote as $\lambda^{S}_{max}$, directly comparable to the isotropic dimensionless EPC strength $\lambda$ reported in literature. The largest $\lambda^{T}_{i}$ ($\lambda^{T}_{max}$) gives the leading odd-parity spin-triplet pairing strength.